# Automted Workload Mapping with 3D Heterogeneous Electronic-Photonic Processing-in-Memory Platform HPiMAP: A Framework for Mapping Hybrid DNNs onto Heterogeneous 3D-Stacked Electronic-Photonic Processing-in-Memory Architectures HPiMAP: A Multi-Objective DNN Mapping Framework on Heterogeneous Electronic-Photonic Processing-in-Memory Architectures

### v1

@@ III-D1 Stage 1: Latency-Energy Pareto Optimization (PO)
In this stage, we perform a multi-objective search that focuses on exploring the Pareto front of energy–latency trade-offs. Since only the number of rows mapped to a tier impacts the inference latency and energy, not specific row indices, we are allowed to significantly reduce the exponential mapping space. The pruned search space now reduces from to . For Pythia-70M, the space is reduced from to 8.61037.
@@
We employ a tailored evolutionary algorithm NSGA-II [19] to solve stage 1, which queries the tier-specific architecture simulators to get the fitness for a particular mapping. The algorithm of stage 1 is detailed in Alg. 1. The simplified energy-latency Pareto optimization problem is
@@
III-D2 Stage 2: Acc-Driven Row-Remap (RR)
In the accuracy optimization stage, we iteratively reassign weight rows from the most noise-prone tier to the tier offering higher accuracy to ensure the process converges to the required accuracy. Although the initial weight mapping (acquired in Stage 1) provides a per-layer distribution, we convert it into a global distribution across all tiers for the entire model. Throughout the reassignment process, we continuously verify that no tier’s memory usage exceeds its capacity. The algorithm finalizes and returns the current assignment once the best tier’s memory is fully occupied or the accuracy requirement is satisfied.
IV Evaluation Results
IV-A Evaluation Setup
Our experiments use the Pythia-70M model [23] trained on the TinyStories dataset [24] as our baseline for studies and analysis. In addition, we evaluate the MobileViT-S model [25] on two distinct datasets: a military asset [26] and a medical chest X-ray dataset [27] to prove our framework’s versatility. All models and datasets are shown in Table II and III.
| Datasets | Dataset Size | Training Set | Test Set | Benchmark |
| TinyStories | 1.92GB | 2.12M rows | 22K rows | 1.1017 (PPL) |
| Military Assets | 4.19 GB | 21978 samples | 1396 samples | 0.8972 (Acc) |
| Chest X-Ray Images | 1.24 GB | 5216 samples | 624 samples | 0.9533 (Acc) |
@@ IV-A Evaluation Setup
| Pythia | 70 M | 24 | 24 | 0 | 6 | 12 |
The models we evaluate retain their original architectures, modified only by including learned step quantization [28]. We first train all models from scratch in an 8-8-8-bit (input–weight–output) configuration, then fine-tune a 6-6-8-bit variant from the 8-bit checkpoint. This fine-tuning step helps maintain a smooth distribution when moving to lower bit precision, accommodating tiers with constrained precision. Note that this fine-tuned 6-6-8-bit model is only used in the RR stage. All experiments are performed on a Linux OS server with NVIDIA A6000 GPUs and AMD EPYC 7763 64-core processors.
IV-B Results
IV-B1 Network-on-Chip Interconnect Cost
Using BookSim NoC simulator [22], we simulate the performance of 2.5D and 3D topologies with inputs of size [8, 3, 32, 32] between two Conv2D layers, shown in Fig. 3. The experiment shows an improvement of 40% and 41% in latency and energy cost, respectively.
IV-B2 Pareto Optimization (PO)
We begin our optimization process by initializing the population and then allowing the NSGA-II [19] algorithm to search for Pareto-optimal solutions of mapping Pythia-70M. Figure 4 illustrates how energy and latency are jointly optimized throughout this evolutionary search. Figure 5 ➊ shows the layer-wise workload distribution among three tiers (Left) and the row-wise mapping results of each layer (Right) after PO.
IV-B3 Accuracy-Driven Row-Remapping (RR)
@@
| 100% SRAM | 10.21 | 13.79 | 1.1017 | 291.92 | 4.70 | 0.953 | 291.92 | 4.70 | 0.897 | ||||||
| 100% ReRAM | 14.73 | 13.44 | 1.1128 | 583.54 | 3.97 | 0.949 | 583.54 | 3.97 | 0.888 | ||||||
| 100% TeMPO | 0.91 | 8.92 | 2.2272 | 2.17 | 15.71 | 0.711 | 2.174 | 15.71 | 0.507 | ||||||
@@
Once we obtain the set of Pareto-optimal solutions, there are inevitably multiple candidates that balance our objectives in different ways. To narrow this down, we explicitly evaluate the accuracy for each of these candidates and select the one that yields the best final accuracy. As a preliminary step, we calculated each row’s sensitivity using a Taylor expansion approach using Eq. (4). Figure 5 right side shows the row-wise sensitivity distribution of each layer sorted from the most sensitive layer to the least one. Leveraging these sensitivity insights, we then gradually shift a portion of the photonic chip’s workload towards SRAM. Since SRAMs are not affected by low-precision and noise-related issues, this rebalancing can help retain the model’s accuracy. As shown in Fig. 6, the PPL steadily improves as the optimization progresses. We impose a 0.1 tolerance relative to the noise-free PPL for RR. Comparing the final workload assignment after RR (Fig. 5 ➋) with the one in PO (Fig. 5 ➊), we observe a noticeable shift in workload from the Photonic tier to SRAM.
| Strategy | Latency (ms) | Energy (mJ) |
Perplexity | LEP Score | ||
| 100% SRAM | 10.21 | 13.79 | 8-8-8 | 1.1017 | 0.5580 | ||
| 100% ReRAM | 14.73 | 13.44 | 8-8-8 | 1.1128 | 0.6428 | ||
| 100% TeMPO | 0.91 | 8.92 | 6-6-8 | 2.2272 | 0.3333 | ||
| Equal Distribution | 4.90 | 12.02 | Mixed | 1.1861 | 0.3339 | ||
[…]
IV-B4 Heterogeneous vs. Homogeneous
@@
To validate our approach, we compare three methods: (1) our proposed framework, (2) an intuitive “equal workload” partition, and (3) a homogeneous workload assignment. Results are shown in Table V. From the table, PO achieves superior efficiency: it reduces latency by 3.66 and energy consumption by 1.22 compared to the equal-workload method, and by 6.42 on latency and 1.22 on energy on average compared to the homogeneous assignment. However, under the PO workload assignment, the model’s performance does not meet the 0.1-difference constraint; after RR, it improves substantially while still maintaining a great latency and energy profile. To facilitate a comprehensive comparison, we introduce a Latency–Energy–Performance (LEP) score, which normalizes and averages three metrics such that lower scores indicate better overall efficiency. As shown in Table V, PO achieves the lowest LEP score; following RR, performance increases more while retaining a second low LEP value. Figure 7 shows each layer’s latency and energy.
IV-B5 Main Results
@@
Lastly, we compare the result of H3PIMAP mapping Pythia-70M (trained on TinyStories) and MobileViT-S (trained on Chest X-Ray and Military Assets) to the proposed 3D heterogeneous architecture with homogeneous mapping, shown in Table IV. Homogeneous mapping shows excessive latency overhead in PIM tiers and unacceptable accuracy degradation in the photonic tier. H3PIMAP mapping results show the solution where the latency and energy are lowered by 3.47 and 2.74, respectively, while maintaining a comparable model accuracy.
V Conclusion
@@
This work presents H3PIMAP, a heterogeneity-aware mapping framework that optimally distributes hybrid DNN workloads across a novel Electronic-Photonic-PIM heterogeneous AI computing architecture. By integrating multi-objective optimization with accuracy-driven remapping, H3PIMAP achieves a Pareto-optimal balance of latency and energy efficiency while ensuring robust model inference against hardware non-ideal variations. On vision and language benchmarks, H3PIMAP achieves a 3.47 reduction in latency and a 2.74 improvement in energy efficiency compared to standard mapping on homogeneous systems. These results highlight the transformative potential of heterogeneous AI hardware, where the synergistic integration of electronic PIM and photonics, coupled with efficient software workload mapping, unlocks new frontiers in next-generation AI acceleration.

### v2

@@ 3.4.1. Stage 1: Latency-Energy Pareto Optimization (PO)
In the first stage of H3PIMAP, we perform a multi-objective search that focuses on exploring the Pareto front of energy–latency trade-offs. Since only the number of matrix rows mapped to a tier impacts the system inference latency and energy, not specific row indices, we are allowed to significantly reduce the search cost in the exponential mapping space. The pruned search space now reduces from combinations to a much smaller space with only combinations. As an example, for Pythia-70M, the search space is reduced from to 8.61037.
@@
We employ a tailored evolutionary algorithm NSGA-II (Deb et al., 2002) to solve stage 1, which queries the tier-specific architecture simulators for the fitness of a particular mapping. The algorithm of stage 1 is detailed in Alg. 1. The simplified energy-latency Pareto optimization is
@@
3.4.2. Stage 2: Acc-Driven Row-Remap (RR)
In the accuracy optimization stage, we iteratively reassign weight matrix rows from the most noise-prone tier to the tier offering higher accuracy to ensure the process converges to the required accuracy. Although the initial weight mapping (acquired in Stage 1) provides a per-layer distribution, we convert it into a global distribution across all tiers for the entire model.
3.5. System Feasibility of H3PIMAP
The H3PIMAP architecture adopts a four-tier 3D-stacked design, with each tier fabricated separately and integrated via standard packaging solutions. This modular approach avoids the yield issues of monolithic integration and is compatible with existing foundry technologies. Its feasibility is supported by increasing academic and industrial demonstrations. Prototypes have stacked SRAM on PICs for near-memory photonic control (Pintus et al., 2025), and bonded photonic-electronic test chips show compatibility with standard 3D stacks without exceeding power, thermal, or yield constraints (Hsia et al., 2021). Researchers are exploring fully optical PIM fabrics that eliminate electronic data movement (Ríos et al., 2019), and industry efforts, such as Lightmatter’s co-packaged optical compute planes with CMOS memory and logic (Long et al., 2019), further validate the practicality of heterogeneous 3D electronic-photonic-PIM integration explored in this emerging study.
| Datasets | Dataset Size | Training Set | Test Set | Benchmark | ||
| TinyStories | 1.92GB | 2.12M rows | 22K rows |
| Military Assets | 4.19 GB | 21978 samples | 1396 samples | 0.8972 (Acc) | ||
| Chest X-Ray | 1.24 GB | 5216 samples | 624 samples | 0.9533 (Acc) |
@@ 3.5. System Feasibility of H3PIMAP
| Pythia-70M | 70 M | 24 | 24 | 0 | 6 | 12 | ||||||||||
| Pythia-2.8B | 2.8 B | 128 | 128 | 0 | 32 | 64 |
4. Evaluation Results
4.1. Evaluation Setup
Our experiments use the Pythia-70M (Biderman et al., 2023) trained on the language task TinyStories dataset (Eldan and Li, 2023) as baseline for analysis. We evaluate the MobileViT-S vision model (Mehta and Rastegari, 2021) on two distinct datasets: a military asset (Rawsi18, 2023) and a medical chest X-ray dataset (Jaeger et al., 2014) to prove our framework’s versatility. All models and datasets are shown in Table 2 and 3.
The models we evaluate retain their original architectures, modified by learned quantization (Esser et al., 2019). We first train all models from scratch in an 8-8-8-bit (input–weight–output) configuration, then fine-tune a 6-6-8-bit variant from the 8-bit checkpoint. This fine-tuning step helps maintain a smooth distribution when moving to lower bit precision, accommodating tiers with constrained precision. Note that this fine-tuned 6-6-8-bit checkpoint is used whenever a workload is evaluated on a configuration that assigns any portion to the photonic tier. All experiments are performed on a Linux OS server with NVIDIA A6000 GPUs and AMD EPYC 7763 64-core processors.
4.2. Results
4.2.1. Network-on-Chip Interconnect Cost
Using BookSim NoC simulator (Jiang et al., 2013), we simulate the performance of 2.5D and 3D interconnect topologies with inputs of size [8, 3, 32, 32] between two Conv2D layers, shown in Fig. 3. The experiment shows an improvement of 40% and 41% in latency and energy cost, respectively.
4.2.2. Pareto Optimization (PO)
We initialize the population and let the NSGA-II (Deb et al., 2002) algorithm seek Pareto-optimal mappings for Pythia-70M. Figure 4 charts energy–latency co-optimization during the evolutionary run, while Fig. 5 ➊ depicts layer-wise workload across three tiers (left) and row-wise mapping per layer (right) after PO.
4.2.3. Accuracy-Driven Row-Remapping (RR)
Once we obtain the set of Pareto-optimal solutions, there are inevitably multiple candidates that balance our objectives in different ways. To narrow this down, we explicitly evaluate the accuracy for each of these candidates and select the one that yields the best final accuracy. As a preliminary step, we calculated each row’s sensitivity using a Taylor expansion approach using Eq. (4). Figure 5 right side shows the row-wise sensitivity distribution of each layer, sorted from the most sensitive layer to the least sensitive one. Leveraging these sensitivity insights, we then gradually shift a portion of the photonic chip’s workload towards SRAM. Since SRAMs are not affected by low-precision and noise-related issues, this rebalancing can help retain the model’s accuracy. In Fig. 6, the PPL steadily improves as the optimization progresses. We impose a 1.0 tolerance relative to the noise-free PPL for RR. Comparing the final workload assignment after RR (Fig. 5 ➋) with the one in PO (Fig. 5 ➊), we observe a noticeable shift in workload from the Photonic tier to SRAM.
| Strategy | Latency(ms) | Energy(mJ) |
Perplexity | LEP Score | ||
| 100% SRAM | 10.21 | 13.79 | 8-8-8 | 20.329 | 1.673 | ||
| 100% ReRAM | 14.73 | 13.44 | 8-8-8 | 20.340 | 1.931 | ||
| 100% TeMPO | 0.91 | 8.92 | 6-6-8 | 23.839 | 1 | ||
| Equal Distribution | 4.90 | 12.02 | Mixed | 22.413 | 1.519 | ||
[…]
@@ 4.2.3. Accuracy-Driven Row-Remapping (RR)
Latency(ms) | Energy(mJ) |
| 100% SRAM | 10.21 | 13.79 | 20.329 | 411.83 | 540.28 | 14.334 | 291.92 | 4.70 | 0.953 | 291.92 | 4.70 | 0.897 | ||||||||
| 100% ReRAM | 14.73 | 13.44 | 20.340 | 581.62 | 529.79 | 14.604 | 583.54 | 3.97 | 0.949 | 583.54 | 3.97 | 0.888 | ||||||||
| 100% TeMPO (Zhang et al., 2024) | 0.91 | 8.92 | 23.839 | 35.86 | 361.92 | 16.419 | 2.17 | 15.71 | 0.883 | 2.17 | 15.71 | 0.827 | ||||||||
@@
4.2.4. Heterogeneous vs. Homogeneous
@@
To validate our approach, we compare three methods: (1) our proposed framework, (2) an intuitive “equal workload” partition, and (3) a homogeneous workload assignment. Results are shown in Table 4. From the table, PO achieves superior efficiency: it reduces latency by 3.66 and energy consumption by 1.22 compared to the equal-workload method, and by 6.42 on latency and 1.22 on energy on average compared to the homogeneous assignment. However, under the PO workload assignment, the model’s performance does not meet the 1.0-difference constraint; after RR, it improves substantially while still maintaining a good latency and energy profile. For a comprehensive comparison, we introduce a Latency–Energy–Performance (LEP) score, defined as the sum of the normalized values of each metric group, where lower scores indicate better efficiency. From Table 4, we see that PO pushes latency and energy close to the pure photonic mapping, and PO+RR then recovers accuracy to deliver the best overall LEP score. Figure 8 shows each layer’s latency and energy after H3PIMAP optimization.
4.2.5. H3PIMAP Effectiveness
| Optimizing Stop Threshold | |||
| Latency ms, Energy mJ, Perplexity | |||
| Strategy | Time Consuming (mins) | ||
84.2 | ||
| H3PIMAP | 51.3 | ||
To demonstrate H3PIMAP ’s effectiveness, we measured optimization time against a full three‑objective evolutionary search, shown in Table 6. A common termination criterion was imposed: latency < 2.5 ms, energy < 10.5 mJ, and perplexity < 1.0, slightly looser than H3PIMAP ’s final solution to avoid bias. Under this criterion, H3PIMAP converged in 51.3 minutes, whereas the pure multi‑objective search required 84.2 minutes, yielding a 39% speed‑up. We also benchmarked our mapping quality against 100 randomly sampled workload assignments, shown in Fig. 7. The first‑stage Pareto Optimization (PO) drives latency and energy deep into the lower‑left corner of the design space; the second‑stage Row Remapping (RR) then recovers model fidelity with only a modest retreat in efficiency. The resulting PO + RR solution dominates most sampled points, achieving substantially lower energy and latency while still meeting the accuracy bound.
4.2.6. Main Results
@@
Lastly, we compare the result of H3PIMAP mapping Pythia-70M (trained on TinyStories), Pythia-2.8B (fine-tuned on TinyStories), and MobileViT-S (trained on Chest X-Ray and Military Assets) to the proposed 3D heterogeneous architecture with homogeneous mapping, shown in Table 5.
Homogeneous mappings (i.e., 100% mapped to a single tier type) either incur large latency in pure PIM tiers or fail to meet the accuracy target on the pure photonic tier. Under matched-quality constraints (NLP models: PPL 1.0; vision models: Acc 0.02), H3PIMAP delivers an average of 3.32 lower end-to-end latency versus the best valid homogeneous baseline across NLP and vision tasks. The advantage is strongest on LLMs: H3PIMAP achieves 77.0 lower latency and 14.6 lower energy on Pythia-2.8B at marginal perplexity discrepancy.
5. Conclusion
@@
This work presents H3PIMAP, a heterogeneity-aware mapping framework that optimally distributes hybrid DNN workloads across a novel Electronic-Photonic-PIM heterogeneous AI computing architecture. By integrating multi-objective optimization with accuracy-driven remapping, H3PIMAP achieves a Pareto-optimal balance of latency and energy efficiency while ensuring robust model inference against hardware non-ideal variations. Across language and vision benchmarks, H3PIMAP delivers a 3.32 reduction in latency compared to homogeneous systems and naïve mappings; on LLMs, it further reduces latency by 77.0 and energy consumption by 14.6 under the constraint of matched task quality. These results highlight the transformative potential of heterogeneous AI hardware, where the synergistic integration of electronic PIM and photonics, coupled with efficient software workload mapping, unlocks new frontiers in next-generation AI acceleration.